### v4

# WorldGUI: An Interactive Benchmark for Desktop GUI Automation from Any Starting Point

Abstract
Recent progress in GUI agents has substantially improved visual grounding, yet robust planning remains challenging, particularly when the environment deviates from a canonical initial state. In real applications, users often invoke assistance mid-workflow, where software may be partially configured, steps may have been executed in different orders, or the interface may differ from its default setup. Such task-state variability is pervasive but insufficiently evaluated in existing GUI benchmarks. To address this gap, we introduce WorldGUI, a benchmark covering ten widely used desktop and web applications with tasks instantiated under diverse, systematically constructed initial states. These variations capture realistic human–computer interaction settings and enable diagnostic evaluation of an agent’s ability to recover, adapt plans, and handle non-default contexts. We further present WorldGUI-Agent, a simple and model-agnostic framework that organizes planning and execution around three critique stages, improving reliability in dynamic environments. Experiments demonstrate that state-of-the-art GUI agents exhibit substantial performance degradation under non-default initial conditions, revealing limited robustness and fragile planning behaviors. Our benchmark and framework provide a foundation for developing more adaptable and reliable GUI agents. The code and data are available at https://github.com/showlab/WorldGUI.
1 Introduction
Table 1: Comparison with other online GUI benchmarks. WorldGUI is a unique benchmark that embraces diverse initial states and better reflects the authentic interactions in GUI scenarios. Env?: Indicates whether an environment is required to be deployed.
Benchmark Software Task Platform Env? Inst. Video? GT Plan Task State
variability?
WebArena (Zhou et al., ) 6 812 Web Yes
VisualWebArena (Koh et al., 2024) 3 910 Web Yes
WebVoyager (He et al., 2024) 15 643 Web Yes
AutoDroid (Wen et al., 2024) 13 158 Android OS Yes
AndroidWorld (Rawles et al., 2024) 20 116 Android OS Yes
AgentStudio (Zheng et al., 2025) 9 205 Desktop + Web Yes
Mobile-Eval (Wang et al., 2024) 10 30 Android OS Yes
APPAgent (Zhang et al., 2023) 10 50 Android OS Yes
OSWorld (Xie et al., 2024) 10 369 Desktop Yes
AssistGUI (Gao et al., 2024) 9 100 Windows No
WindowAgentArena (Bonatti et al., 2024) 11 154 Windows Yes
WorldGUI 10 611 Win. + Web No
2 Introduction
Graphical User Interface (GUI) automation has emerged as a promising direction for productivity-oriented AI systems. With the rapid advancement of Multimodal Large Language Models (MLLMs) such as GPT-5 (Singh et al., 2025) and Claude-4-Sonnet (Anthropic, 2025), GUI agents have demonstrated the potential to assist users in software usage, file management, office design, coding, and web browsing, reducing repetitive work and improving workflow efficiency.
Unlike traditional computer vision tasks such as image recognition (He et al., 2016) or visual question answering (Antol et al., 2015; Goyal et al., 2017), GUI automation operates in a highly dynamic environment where the interface state evolves as a result of user actions, system configurations, or partial task progress. However, existing online GUI benchmarks, including WebArena (Zhou et al., ), WebVoyager (He et al., 2024), OSWorld (Xie et al., 2024), AssistGUI (Gao et al., 2024), and WindowsAgentArena (Bonatti et al., 2024), primarily evaluate agents from a single, canonical initial state and measure success only at the end of the trajectory. Such static evaluation overlooks several common situations in real-world usage: (1) software interfaces are often not in their default configurations, (2) users frequently request assistance from intermediate states of partially completed tasks, and (3) agents with the same low success rate (e.g., 20%) may differ substantially in their robustness, planning stability, and self-correction behavior differences that cannot be revealed under a single start-state protocol.
In parallel, robustness-oriented GUI benchmarks such as GUI-Robust (Yang et al., 2025) examine how agents respond to interface-level perturbations, including layout shifts, appearance variations, and transient pop-ups. These benchmarks vary the visual presentation while keeping the task goal and initial state fixed, providing valuable insights into perception robustness. In contrast, WorldGUI does not alter UI skin, resolution, or layout; instead, it targets task-progress–level variation by modifying the executed steps themselves (Add-, Trim-, and Adjust-step). This form of task-state variability, such as missing, redundant, or reordered steps, poses challenges fundamentally different from interface perturbations, making the two robustness dimensions complementary and largely orthogonal.
To complement existing benchmarks, we focus on robustness to intermediate starting states, a ubiquitous but under-explored aspect of GUI interactions. We introduce WorldGUI, a benchmark that systematically constructs diverse initial states for each task by applying controlled pre-actions. These pre-actions modify task progress itself (e.g., steps partially completed, over-executed, or executed in a different order), allowing tasks with the same goal to begin from multiple realistic contexts. This design mirrors how real users summon assistance in the middle of workflows and enables diagnostic evaluation of an agent’s ability to recover, adapt plans, and handle contextual variability. WorldGUI covers 10 widely used desktop applications and provides 611 tasks, each paired with a user query, instructional video, and project file. To construct meaningful state variations, trained annotators first demonstrate ground-truth (GT) plans, after which we generate augmented initial states via pre-actions.
In addition, we propose WorldGUI-Agent, a simple and general GUI agent framework grounded in critical-thinking principles, an aspect that has received relatively limited attention in prior GUI agents (Hong et al., 2024; Cheng et al., 2024; Lai et al., 2024; Agashe et al., 2025a; Wu et al., 2024). Dynamic GUI environments often deviate from expected states due to user behavior or system configurations, making it essential for agents to detect mismatches and adapt their plans. Through an analysis of real-world GUI scenarios, we distill three design principles that we argue are fundamental for GUI automation: (1) Post-Planning Critique, (2) Pre-Action Validation, and (3) Post-Action Evaluation. These components enable WorldGUI-Agent to refine plans, verify step prerequisites, and evaluate action outcomes, thereby improving reliability in dynamic settings.
Experiments reveal three key observations: (1) a substantial gap remains between humans and current agents, with models such as UI-Tars and GPT-5.1 still struggling on our curated desktop GUI tasks; (2) performance degrades sharply under non-default initial states, highlighting limited robustness and fragile planning capabilities; and (3) WorldGUI-Agent significantly improves robustness and overall task success across WorldGUI and WindowsAgentArena, even when built upon weaker base models and without prompt tuning, demonstrating the importance of multi-level critique for dynamic GUI automation (see Section 6).
3 Related Work
Figure 1: WorldGUI. Left: WorldGUI creates pre-actions for each meta task, leading to different initial states. It successfully reflects the real-world human-computer interaction process. Right: components in WorldGUI.
3.1 GUI Benchmarks
GUI benchmarks provide the foundation for evaluating computer-use agents across web, desktop, and mobile environments. Early interactive benchmarks such as WebShop (Yao et al., 2022), WebArena (Zhou et al., ), and WebVoyager (He et al., 2024) center on browser tasks with long-horizon goals. OSWorld (Xie et al., 2024) and mobile benchmarks, including MobileAgent (Wang et al., 2024) and AppAgent (Zhang et al., 2023) broaden coverage to multi-application desktop and mobile ecosystems. AssistGUI (Gao et al., 2024) and WindowsAgentArena (Bonatti et al., 2024) provide online execution environments for Windows-based tasks.
Recent benchmarks emphasize scalability and online evaluation. UIExplorer (Nica et al., 2025) offers a large-scale environment for GUI exploration in both structured-DOM and pure screen modes. GUI-World (Chen et al., 2025) provides multi-application desktop tasks with online execution and standardized initial states. Robustness-oriented suites such as GUI-Robust (Yang et al., 2025) examine agents under interface-level perturbations (e.g., layout shifts, visual noise, pop-ups), focusing on perception and action robustness. While these benchmarks significantly advance GUI automation, they overwhelmingly assume a single canonical initial state for each task. They do not explicitly evaluate task-state variability, such as partially completed workflows, missing or redundant steps, or alternative intermediate contexts common in real user workflows. Our benchmark, WorldGUI, complements existing online environments by systematically constructing diverse initial states for the same task goal via controlled pre-actions, enabling diagnostic evaluation of robustness, recoverability, and planning stability in dynamic GUI environments.
3.2 GUI Agents
Recent progress in GUI agents spans perception, grounding, planning, and full computer-use control. CogAgent (Hong et al., 2024) improves GUI perception, while SeeClick (Cheng et al., 2024) and SeeAct (Zheng et al., 2024) focus on grounding UI elements for more reliable action prediction. MobileAgent (Wang et al., 2024) and AppAgent (Zhang et al., 2023) extend these capabilities to mobile environments. On desktop and OS-level tasks, AssistGUI (Gao et al., 2024) adopts a Plan-Act pipeline with a single post-action checker. Agent-S (Agashe et al., 2025a) introduces experience-augmented hierarchical planning with a Manager-Worker architecture, a self-evaluator, and an Agent-Computer Interface (ACI). Agent-S2 (Agashe et al., 2025b) further incorporates specialist grounding (Mixture-of-Grounding) and Proactive Hierarchical Planning, achieving strong performance across OSWorld, WindowsAgentArena, and AndroidWorld. Beyond inference-time frameworks, several works train GUI agents or their components. Aguvis (Xu et al., 2024) and UGround (Gou et al., 2025) train pure-vision or universal grounding models at scale. AgentTrek (Xu et al., 2025) synthesizes multi-step trajectories from web tutorials for supervised training, and SE-GUI (Yuan et al., 2025) applies reinforcement learning to strengthen grounding robustness. While these approaches enhance perception, grounding, or hierarchical planning, they typically employ only a single verification stage (e.g., self-evaluation or post-action checking). In contrast, our framework WorldGUI-Agent focuses on structuring the verification process itself for dynamic GUI environments. WorldGUI-Agent introduces three complementary critique stages: Post-Planning Critique, Pre-Action Validation, and Post-Action Evaluation, enabling agents to detect state mismatches, assess step readiness, and repair execution failures. This multi-level critique design is orthogonal to hierarchical planners and training-based pipelines, and can be combined with them without additional model training. A component-wise comparison is provided in Table 3.
4 WorldGUI Benchmark
Following prior work (Xie et al., 2024; Zhou et al., ), we model GUI automation as a partially observable Markov decision process (POMDP). A complete specification of the POMDP components and the execution environment is provided in Appendix A.
4.1 Data Source
WorldGUI consists of a broad spectrum of widely-used desktop applications, which can be categorized into five main groups: (i) Office, includes PowerPoint, Word, Excel, and Adobe Acrobat; (ii) Windows Usage, includes System Settings and File Management; (iii) Web Usage, includes the configuration of Youtube and website operations; (iv) Coding, focus on the customization, configuration and editing of Visual Studio Code (VSCode); (v) Media, operating VLC player for video editing and creation.
Table 2: Task category, task activities, and project file of the desktop applications in WorldGUI.
Category Applications All Task Avg. Length Task Activities Project File Type
Office PowerPoint 64 7.1 Change the content style and layout; Design new effects project.pptx
Office Word 63 4.1 Formatting the content style and layout project.docx
Office Excel 70 5.1 Table formatting; Data management and processing project.xlsx
Office Adobe Acrobat 66 5.1 Automatic add electric signature; Document management project.pdf
Coding VSCode 56 4.4 Code editing; Software configuration vscode.exe
Windows Usage Settings 69 7.1 Advanced personalized and safety settings; ms-Settings
Windows Usage File Explorer 44 5.1 File management: Add, delete, rename, and move files explorer.exe
Web Usage Web Browser 59 7.8 Web operation web browser + URL
Web Usage Youtube (Online) 61 4.8 Video and account configurations web browser + URL
Media VLC Player 59 10.8 Video editing and creation project.mp4
Total (Average) 611 6.0
Figure 2: The left shows software taxonomy in WorldGUI. The right shows the distribution of task length.
4.2 Pipeline of Data Construction
We engaged four annotators and developed the necessary scripts to structure and format the data. To reduce annotator-style variance, all annotators followed a standardized annotation protocol. Additionally, to facilitate ground truth (GT) plan generation and pre-action generation, we implemented simple agent-based methods to collect the relevant data. The overall data construction pipeline comprises six steps, as detailed below.
Raw Video Collection. We collect raw videos from the YouTube website as there are a lot of high-quality tutorials for desktop applications with high views. For each software, we ask the annotators to watch the videos first and download them via the diversity of software usage.
Instruction Video Preparation. After obtaining the raw videos, we write the script codes to cut the lengthy and noisy videos into the sub-clips (30 seconds to 3 minutes) that serve as the instructional video.
User Query Generation. After obtaining the instructional videos, annotators are asked to manually write user queries corresponding to each video. For example, a user query for a task involving File Explorer might be: “Please compress the project.mp4 into an MPEG-4 file optimized in 1080p.”
Project File Preparation. Following the AssistGUI (Gao et al., 2024), we create the project file for each task to ensure reproducibility without relying on resource-intensive virtual machines (Xie et al., 2024) or Docker environments (Bonatti et al., 2024). This approach guarantees that the testing process begins from a consistent state. When combined with pre-actions, it enables augmentation of the same task with various initial states.
GT Plan Generation. Given a user query $q$ and instructional video $v$, we automatically generate an initial draft plan using a GPT-4o–based agent. Annotators then watch the video and execute the task while reviewing the draft plan, correcting any inaccurate or missing steps to obtain the final GT plan. GT plans are used solely for constructing pre-actions and are never provided to or used by any evaluated agent.
Pre-Actions Generation. To vary the task, we propose introducing pre-actions before the task begins. These pre-actions are created by annotators and involve corresponding scripts and agents. They are written in Python code, for example: from pyautogui import click, $\n rightClick$(800,400). The pre-actions primarily serve two purposes: 1) Simulating Intermediate Task States: Pre-actions can complete specific steps of a task, creating a starting point from an intermediate state. This approach addresses scenarios where users may invoke GUI assistant at any time. For example, if the task involves opening a dropdown menu, the pre-action may pre-open the menu. If the agent fails to recognize this precondition and follows its plan to click the menu again, it might inadvertently close the menu, causing task failure. 2) Introducing Diverse Initial Context States: Pre-actions can also introduce variations in the initial state, such as opening random tabs or settings. This ensures that the starting state is unconventional, challenging the agent to adapt by modifying its plan or adding necessary new steps. See example in Figure 5. Each pre-action was manually verified to ensure the task remained solvable.
4.3 Data Statistics
Overall. We present the detailed task taxonomy and lengths about WorldGUI in Figure 2 and Table 2. A total of 111 meta tasks were collected from these applications, with each task being augmented  5 times based on the task’s functionality, resulting in 500 augmented tasks. In total, WorldGUI comprises 611 tasks, and every task has almost 6 variation instances, which is capable of reflecting the real-world interactions of the GUI environment. All data are available for reference. [^1] https://huggingface.co/datasets/anonymousABC/WorldGUI-Bench
Figure 3: The distribution of different augmentation types.
Figure 4: The distribution of different task difficulty.
Figure 5: An example of augmenting one GUI task with manually aug the initial state and then using the execution scripts and corresponding agents to obtain the pre-action for each augmented case.
Augmentation tasks type analysis. As we summarize, the real GUI scenarios include: (1) The software interface is not in its default state. (2) The human-computer interactions may start from the intermediate state of a specific task. Our augmentations lie in two main groups: (1) simulating the intermediate states and (2) introducing diverse initial states. We divide the two groups into three types: Add-step, Trim-step, and Adjust-step. For Add-step, it represents various unrelated state augmentations to simulate the scenario that we may start the task in another unrelated task or interfaces, the agent should replan the task to add necessary steps. For Trim-step, it represents that we finish several steps of a long task and make the task in an intermediate state. For Adjust-step, it is usually a small modification of the existing state, such as changing the interface by clicking another Tab or clicking a button to open an unrelated dropdown menu. Most of the time, it would not require new steps to return to the target task procedure. This augmentation may mislead the agent in state understanding, making them jump or miss the key steps. As shown in Figure 3, the manually created augmentations mainly belong to the add-step. Adjust-step could be the second-largest part. Due to the low complexity of the interfaces of File Explorer, we cannot create many augmentations for adjust-step.
Task difficulty analysis. Figure 4 shows the distribution of the task difficulty across desktop applications. We annotated the task difficulty level based the subjective software usage experience. The results indicate that the tasks in Adobe Acrobat and VLC player are more challenging. The tasks in Excel, PowerPoint, and Word are more at the medium and simple levels. By considering the Success Rate and task length on these tasks, one can know that the tasks are easy for humans, but hard for current GUI agents. Overall, the task difficulty of our created data is diverse across different applications.
5 WorldGUI-Agent
Table 3: Comparison with closely related GUI agents.
Method Post-Plan Pre-Action Post-Action
Mobile-Agent / V2
AssistGUI
Agent-S / S2
Mobile-Agent-E
WorldGUI-Agent (ours)
To study robustness under the dynamically perturbed initial states of WorldGUI, we adopt a lightweight, non-training agent framework. As summarized in Table 3, most existing agents rely on a single post-action evaluation module. Rather than proposing a new algorithm, our aim is simply to structure the standard perception–planning–execution loop into clearer verification stages.WorldGUI-Agent introduces three small modules at different abstraction levels: Planner-Critic refines the draft plan; Step-Check uses the screenshot and metadata to decide whether a plan step should be executed, skipped, or rewritten; Actor-Critic validates each action using the current observation and recent action history. Compared to prior self-reflection approaches (e.g., Reflexion, CRITIC), which apply a single global critique, our checks are lightweight and distributed across plan, step, and action levels to better handle task-progress variation. Full module specifications and I/O formats are included in Appendix B.
6 Experimental Results
Table 4: Success rate (%) of different agents on WorldGUI. Human denotes the average performance of four expert participants who have watched the instructional video only once, similar to the model. Meta represents the meta task, while Aug. represents the augmented task.
Method Office Win. Usage Web Coding Media Overall
Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
Plan-Act w/ Gemini-2.0 8.9 3.2 8.3 3.4 28.6 16.2 18.2 2.2 10.0 2.0 6.9
Plan-Act w/ GPT-4o 13.3 10.1 8.3 2.3 23.8 11.1 9.1 2.2 10.0 2.0 8.5
AssistGUI w/ GPT-4o 26.7 16.1 29.2 7.9 33.3 20.2 27.3 11.1 10.0 8.2 16.5
CCU w/ Claude-3.5-Sonnet 28.9 19.3 29.2 14.6 71.4 32.3 54.6 22.2 30.0 6.1 23.6
UI-TARS-1.5 28.9 16.1 12.5 2.2 28.6 9.1 36.7 6.7 0.0 0.0 12.3
Agent S2 33.3 16.5 70.8 59.6 52.4 45.5 45.5 37.8 20.0 16.3 34.2
WorldGUI-Agent (Ours)
w / Gemini-2.0 31.1 17.0 20.8 9.0 38.1 29.3 36.4 11.1 20.0 10.2 19.1
w / GPT-4o 42.2 24.3 41.7 11.2 47.6 35.4 45.5 15.6 40.0 12.2 26.0
w / Claude-3.5-Sonnet 57.8 32.6 50.0 19.1 76.2 46.5 54.6 26.7 50.0 18.4 36.0
w / GPT-5.1 64.4 37.6 62.5 40.4 85.7 53.5 63.6 46.6 60.0 26.5 45.8
Human 88.9 83.5 100.0 89.9 95.2 80.8 81.8 77.8 90.0 85.7 85.3
Table 5: Success rate (%) of our WorldGUI-Agent (w/ GPT-4o) with the ablation of different critical modules.
Method Office Win. Usage Web Coding Media Overall
Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
Full Model 42.2 24.3 41.7 11.2 47.6 35.4 45.5 15.6 40.0 12.2 26.0
– w/o Planner-Critic 31.1 17.0 20.8 9.0 38.1 25.3 36.4 11.1 20.0 10.2 18.5
– w/o Step-Check 31.1 19.3 20.8 9.0 33.3 28.3 45.5 13.3 20.0 8.2 19.8
– w/o Actor-Critic 15.6 7.8 4.2 3.4 28.6 17.2 0.0 8.9 10.0 6.1 9.7
Table 6: Success rate (%) of our WorldGUI-Agent with the ablation of Instructional Video.
Method PPT Word Excel Acrobat VSCode Overall
Full Model 45.5 36.4 50.0 36.4 45.5 42.9
w/o Inst. Video 45.5 27.3 25.3 18.2 27.3 28.6
Table 7: Running time comparison.
Method Executed Steps Time (seconds)
Agent-S 10 131.98
Agent-S2 9 108.65
WorldGUI-Agent 24 129.55
Table 8: Task counts and success rates (SR, %) across different augmentation types for WorldGUI-Agent (GPT-5.1).
Augmentation Type #Tasks SR (%) #Success
Add-step 255 34.1 87
Trim-step 119 52.1 62
Adjust-step 130 43.1 56
Total 500 41.0 205
[…]
Evaluation. Given that our WorldGUI includes 611 GUI tasks, we engaged four participants with strong coding and software backgrounds to test all tasks and document their evaluation results. Metric. Following the previous works, we use Success Rate (SR) as the metric.
Baselines. We implement the baseline approach called Plan-Act with different MLLMs as the base model. It focuses on investigating the basic capabilities of task planning and action prediction. Additionally, we compare our WorldGUI-Agent with two agentic frameworks and two SOTA GUI models: AssistGUI (Gao et al., 2024), Agent-S2 (Agashe et al., 2025b), Computer Use (Claud-3.5-Sonnet) (Anthropic, 2024), and UI-TARS-1.5 (Qin et al., 2025). AssistGUI and Agent-S2 are two prominent agentic frameworks designed for Desktop GUI Automation, which can plan the task and then execute the task step by step by following the query. We increase the base model to GPT-4o for AssistGUI and Claude-Sonnet-4 of Agent-S2 for fair performance. Claude Computer Use (CCU) is the leading proprietary model specially designed for computer use. We use the open-source implementation OOTB (Hu et al., 2024) as the codebase and then add the subtitle of instructional videos into the input prompt for a fair comparison. We also implement our WorldGUI-Agent with four different MLLMs to illustrate the effectiveness of our proposed universal agent framework.
6.1 Main Results on WorldGUI
Overall. Table 4 reports the success rates (SR) of different agents and human experts on our WorldGUI benchmark, broken down by task type (Meta vs. Aug.) across five categories: Office, Win. Usage, Web, Coding, and Media. From these results we draw the following main conclusions.
A large gap remains between agents and humans. The best-performing agent (WorldGUI-Agent with GPT-5.1) achieves an overall SR of only 45.8%, which is less than half of the 85.3% attained by human experts. This stark contrast underscores the difficulty of our tasks and the need for further advances in desktop GUI automation.
Agents generalize poorly to augmented tasks. Across all methods, performance on Augmentation tasks (which introduce interface or context variations) is substantially lower than on their corresponding Meta tasks. For example, Claude-3.5-Sonnet in the Win. Usage category attains 50.0% on Meta tasks but drops to just 19.1% on Aug. tasks. This highlights the importance to evaluate the agents on various initial conditions.
Desktop GUI tasks pose a greater challenge than web tasks. Every agent scores higher on Web tasks than on desktop application tasks. WorldGUI-Agent with GPT-5.1, for instance, jumps from 85.7% on Web Meta to only 62.5% on Win. Usage Meta, and the gap widens on their Augmentation counterparts. Thus, desktop GUI tasks remain challenges than widely-studied web GUI tasks (Cheng et al., 2024; He et al., 2024).
WorldGUI-Agent consistently outperforms a naive Plan-Act counterpart. By incorporating our three critical modules into the planning and execution loop, WorldGUI-Agent substantially improves success rates over the basic Plan-Act counterpart. Relative to Plan-Act, WorldGUI-Agent raises overall SR by +12.2% with Gemini-2.0, +17.5% with GPT-4o, and +12.4% with Claude-3.5-Sonnet, demonstrating the effectiveness of our critical modules proposed for WorldGUI-Agent.
Specialized GUI models transfer poorly to WorldGUI. Despite achieving strong results on several existing GUI benchmarks, specialized models do not perform well in our setting. For example, UI-TARS-1.5 attains only 12.3% overall SR on WorldGUI, which is notably lower than general-purpose MLLMs when used within our framework (e.g., 36.0% for WorldGUI-Agent with Claude-3.5-Sonnet and 45.8% with GPT-5.1), and even below the CCU baseline built on Claude-3.5-Sonnet (23.6%). Similarly, Agent-S2 reaches 34.2% overall SR, still far from human performance and trailing behind WorldGUI-Agent with GPT-5.1. These results indicate that training pipelines and architectures tuned for single-start or less diverse benchmarks do not automatically transfer to the multi-start, dynamically perturbed settings of WorldGUI. In our evaluation, specialized GUI models exhibit strong grounding but substantially weaker planning, which leads to low success rates even when provided with additional signals such as instructional-video subtitles.
Table 9: Experimental results on WindowsAgentArena. The reported results are from the (Bonatti et al., 2024) and (Agashe et al., 2025b).
Method Office Web Win. System Coding Media Win. Utils Overall
Phi3-V (Bonatti et al., 2024) 0.0 6.9 8.3 0.0 6.2 0.0 3.5
GPT-4o-mini (Bonatti et al., 2024) 0.0 14.9 8.3 0.0 0.0 0.0 4.2
GPT-4o (Bonatti et al., 2024) 0.0 13.7 29.2 0.0 10.3 0.0 8.6
NAVI (Bonatti et al., 2024) 0.0 27.3 33.3 27.3 30.3 8.3 19.5
Agent S (Agashe et al., 2025a) w/ GPT-4o 0.0 13.3 45.8 29.2 19.1 22.2 18.2
Agent S2 (Agashe et al., 2025b) w/ Claude-3.7-Sonnet 7.0 16.4 54.2 62.5 28.6 33.3 29.8
WorldGUI-Agent w/ Claude-3.5-Sonnet 7.0 53.3 45.8 33.3 28.6 33.3 31.2
6.2 Ablation Study
Impact of different critical modules. Table 5 reports an ablation study of WorldGUI-Agent across five application categories. The full model achieves an overall success rate (SR) of 26.0%. Removing any core component leads to notable degradation. Excluding the Planner-Critic reduces SR to 18.5% (–7.5%), highlighting its role in plan refinement. Removing Step-Check lowers SR to 19.8% (–6.2%), mainly affecting multi-step interaction scenarios, suggesting its effectiveness in intermediate error correction. The most severe impact comes from removing the Actor-Critic, which collapses SR to 9.7% (–16.3%), nearly eliminating performance on Coding and Win. Usage tasks, underscoring the necessity of action-level reward feedback. Overall, the three modules provide complementary benefits, enabling robust agent performance.
Impact of Instructional Video. In Table 6, we study the impact of removing the instructional video (subtitles) by modifying the prompt to include only the user query for generating the initial plan. In the Excel applications, we observe a significant performance decline, as their tasks are complex and difficult, and rely more heavily on additional domain knowledge for successful planning. In contrast, the agents performs relatively well on Win. Usage tasks, such as Settings and File Management, are where it has more inherent pretrained prior. These findings underscore the necessity of instructional videos for complex tasks like visual effect design, mirroring how users learning to build a slide often rely on tutorial videos. It is note that it is a setup proposed in our benchmark; one can also use our data without instructional videos.
6.3 Results by Augmentation Type
Across the 500 augmented tasks, WorldGUI-Agent (GPT-5.1) solves 205 cases. As shown in Table 8, Add-step tasks are the most challenging (34.1% SR), since inserted steps create larger state mismatches. Adjust-step tasks show moderate difficulty (43.1%), reflecting the need for fine-grained manipulations. Trim-step tasks are relatively easier (52.1%), as removing redundant steps keeps the interface closer to the meta configuration. These trends highlight how different forms of task-progress variation distinctively impact agent robustness.
6.4 Failure Analysis and Robustness Dimensions
To better understand the limitations of current GUI agents under the dynamic initial states of WorldGUI, we analyze representative failure modes drawn from our qualitative study (Appendix I). We observe three major robustness dimensions that frequently lead to failures.
(1) State-mismatch robustness. When pre-actions introduce interface configurations different from the default view (e.g., an expanded settings dropdown or partially filled input fields), agents often misinterpret the visible state or fail to recognize that previous steps have already been executed. As illustrated in Figure 24 (left), once a submenu overlays other elements, the agent struggles to recover the intended target (e.g., the “System” button), leading to incorrect plan continuation or premature termination.
(2) Fine-grained manipulation robustness. Tasks requiring precise adjustments—such as dragging sliders or manipulating bars to reach a specific value—remain challenging. Even when grounding is correct, the agent may overshoot, undershoot, or repeatedly click without achieving the required adjustment, as shown in Figure 24 (right). This reflects a broader difficulty in performing continuous or multi-step manipulations that depend on spatial granularity.
(3) Ambiguous-visual-context robustness. In the absence of clear text cues, agents exhibit uncertainty in selecting visually similar elements. Figure 25 (left) illustrates a case where the model must choose a centered icon from a symmetric layout, yet struggles due to limited visual discrimination without textual anchors. Similarly, when multiple text elements appear close together, GUI parsers may surface misleading anchors (e.g., the label “Replace with”), causing the agent to click an irrelevant surface-level text region rather than the actual input box (Figure 25, right).
Overall, these failure modes echo our quantitative findings: while grounding on static interfaces is often strong, robustness to state changes, fine-grained manipulations, and text-sparse visual contexts remains limited. These observations highlight the need for future progress in both perception and planning components, particularly under the multi-start, dynamically perturbed conditions introduced by WorldGUI.
6.5 Computational Costs Analysis
We select two SoTA non-training agent frameworks Agent-S (Agashe et al., 2025a) and Agent-S2 (Agashe et al., 2025b) for comparing the computational costs. Take a Windows setting task as an example, we tested on the same desktop PC with an AMD Ryzen 7 5800H CPU. As shown in Table 7, our WorldGUI-Agent shows a competitive running time of 129.55s, as compared with Agent-S (131.98s) and Agent-S2 (108.64s). The main computational costs of our designed modules are largely affected by calling base model. Since the main problem of desktop GUI automation is still the suboptimal performance as shown in Table 4, such computational costs are currently acceptable. There remains a clear tradeoff between performance and time costs in GUI automation, and this challenge is shared across the community.
6.6 Results on WindowsAgentArena
Settings. The official WindowsAgentArena (WAA) evaluation runs inside a Linux-based VM using a pre-built image and a JSON configuration file. Since the essential logic of WAA lies in its JSON specification of the environment state and task instruction, we reconstruct the same environment on a Windows machine by following the provided configuration (e.g., opening specified tabs or launching required applications). We then use the instruction field from the JSON file as the prompt for our WorldGUI-Agent, which executes the corresponding GUI actions on Windows. After completion, we follow WAA’s evaluation protocol to determine task success. When an official evaluation script is available, we apply it directly; otherwise, we perform manual judgment based on the given instructions. We do not optimize the prompts, just apply our agent directly on its tasks. Main results. Table 9 compares WorldGUI-Agent against six agents on the WAA benchmark. WorldGUI-Agent achieves a 31.2% overall SR surpass two SOTA api-calling agents, Agent-S and Agent-S2, even with a weaker base mode,l Claude-3.5-Sonnet. These results underscore the effectiveness of our designed critical modules.
Conclusion
We introduced WorldGUI, a benchmark that evaluates GUI agents under diverse, dynamically perturbed initial states, an aspect largely absent from existing evaluations. By constructing task-progress variations via pre-actions, WorldGUI enables systematic assessment of robustness and recovery in realistic GUI scenarios. We also presented WorldGUI-Agent, a lightweight, universal framework that inserts three verification stages into the standard planning loop to improve stability in dynamic environments. Our experiments show that current agents, including strong commercial models, still struggle under non-default states, while WorldGUI-Agent offers consistent gains.
Impact Statement
This paper introduces a benchmark and agent framework intended to advance research on dynamic GUI automation. Our goal is to improve the evaluation and reliability of GUI agents in controlled environments, which may support future progress in productivity tools, accessibility assistance, and educational software. As with many advances in machine learning and automation, there are potential risks, including misuse of automated GUI agents for large-scale or unauthorized interactions with software systems. Our benchmark does not involve personal or sensitive data and is designed strictly for research use. We encourage practitioners to incorporate appropriate safeguards, such as permission checks, sandboxing, and oversight, when deploying GUI agents in real-world applications. Overall, we believe the broader societal impacts of this work align with those commonly associated with progress in machine learning, and we do not identify specific concerns beyond these general considerations.
Appendix A Task Formulation
GUI Automation. Following prior work (Xie et al., 2024; Zhou et al., ), we model GUI automation as a partially observable Markov decision process (POMDP) $(\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{T},\mathcal{R})$. Given a natural-language query $q$ and optional instructional video $v$, the agent receives an observation $o_{t}$ of the current state $s_{t}$ and outputs an action $a_{t}\in\mathcal{A}$, which updates the environment to $s_{t+1}$. The episode terminates when the task succeeds or fails, and $\mathcal{R}$ returns a binary completion signal.
WorldGUI Tasks. As illustrated in Figure 1, each task in WorldGUI is instantiated under multiple initial states that all converge to the same goal. We create these variants through pre-actions—short executable sequences that shift task progress without altering the UI skin, resolution, or layout. This design enables controlled task-progress perturbations (Add/Trim/Adjust-step) and complements existing online GUI benchmarks; key differences are summarized in Table 1.
Observation Space. The agent observes (i) application metadata $m_{t}$ (e.g., panel structure and parsed UI elements), and (ii) a screenshot $V_{t}$ of the current GUI state, following  (Gao et al., 2024). Metadata provides structured grounding anchors, while screenshots offer holistic visual context for planning and action generation.
Action Space. Actions include raw mouse and keyboard operations (click, drag, type, shortcuts), with mouse actions parameterized by pixel coordinates in $V_{t}$. We use PyAutoGUI [^4] https://pyautogui.readthedocs.io for execution, and represent each action in the form action_type(arguments) (Table 10).
Table 10: The action types and examples.
Action Type Example
Mouse Movement moveTo(120, 200)
Mouse Clicks click(200, 300)
Keyboard Type write(’classes’)
Hotkey hotkey(’ctrl’, ’a’)
Scrolling scroll(-100)
Drag dragTo(120, 220, 2)
Mouse Down and Up mouseDown(); mouseUp()
Press Keys press(’delete’)
Key Down and Up keyDown(’shift’)
Appendix B WorldGUI-Agent: Thinking before Doing
Figure 6: WorldGUI-Agent. The Planner module receives the user query and an instructional video as input and generates an initial plan. This plan is then refined and executed step by step. Before each step is passed to the Actor module, it undergoes a Step-Check. After the Actor produces an action, the Actor-Critic module iteratively verifies the completion of the action and makes corrections.
In this section, we introduce an universal GUI framework WorldGUI-Agent with a core and essential designing principle: critical thinking, which is vital for designing GUI agents capable of handling dynamic environments that have been overlooked in prior GUI agents  (Hong et al., 2024; Cheng et al., 2024; Lin et al., 2024; Zhang et al., 2023; Agashe et al., 2025a). The WorldGUI-Agent includes the five fundamental but essential components as in Figure 6 and an Interaction reasoning loop detailed in Algorithm 1. We summarize our critical designs in the following:
Figure 7: State-Aware Planner and Planner-Critic. The Planner generates an initial plan. Then, the Planner-Critic provides necessary corrections.
$\bullet$ Post-Planning Critique: After the planning phase, a critique module verifies and, if necessary, self-corrects the generated plans to ensure their accuracy.
$\bullet$ Pre-Action Validation: Before executing each subtask, a validation module determines whether the subtask should be executed. This step is crucial, as the current GUI environment may indicate that the subtask is unnecessary or requires modification to align with the current state.
$\bullet$ Post-Action Evaluation: After each action execution, a mechanism evaluates whether the action was successfully completed before proceeding to the next subtask.
These critique designs ensure the reliability and adaptability of WorldGUI-Agent in complex GUI environments.
B.1 State-Aware Planner
The State-Aware Planner processes the instructional video $v$ and the user query $q$ generates an initial plan as shown in the left of Figure 7. We use the speech recognition model Whisper (Radford et al., 2023) to translate the video $v$ into the subtitle and then send it to the MLLM for task planning. The task plan is hierarchically structured as $p=[p_{1},p_{2},...,p_{N}]$ where $p_{i}$ is a text string describing the $i$-th milestone of the task. Under each $p_{i}$, there is a list of subtasks $S^{i}_{1},S^{i}_{2},S^{i}_{N}]$, where $S^{i}_{j}$ is the $j$-th subtask in the $i$-th milestone. To ensure the produced plans fit the GUI environment, we propose incorporating an initial screenshot $V_{0}$ to represent the current state. This additional context allows the agent to output plans that align with the actual state. For example, if the instructional video suggests clicking on the “Layout” tab in the Word application, but the current state (as indicated by the screenshot) shows that the “Layout” tab is already selected, there is no need to perform this action again. By utilizing the visual information from the screenshot, the State-Aware Planner can modify the plans accordingly, rather than strictly following the guidance in the instructional video or the existing knowledge from backbone MLLMs. It also avoids the occlusion issue when not seeing the screenshot.
B.2 Planner-Critic
Post-Planning Critique. The goal of the Planner-Critic is to assess the correctness of the initial plans generated by the State-Aware Planner and provide corrections if needed. This module is designed to ensure the accuracy of the plans while leveraging the self-reflection capabilities of MLLMs. As illustrated in Figure 7, for each Initial Plan, the output consists of four components:
(1) <Flag>: Indicates whether the Initial Plan is correct.
(2) <Feedback>: Identifies the error type, categorized into one of three options: “Wrong Steps,” “Missing Steps,” or “Redundant Steps.”
(3) <Correction>: Provide the corrected plans if the Flag indicates that the Initial Plan is incorrect.
(4) <Reason>: In addition to giving the corrected plans, we force the model to give the reasons. As previous studies (e.g., CoT (Wei et al., 2022), Deepseek-R1 (DeepSeek-AI et al., 2025)) demonstrate that generating reasoning steps along with the answer would enhance the performance.
B.3 Step-Check
Figure 8: Step-Check. This module first checks the step completion status via an MLLM and then navigates to the current task processing.
Pre-Action Validation. After the plan assessment, a navigation mechanism is crucial before sending each subtask $S_{t}=S_{j}^{i}$ at the time step $t$ to the Actor module. To address this, we designed a new module called Step-Check. Through extensive investigation, we discovered that during GUI task testing, perfect execution plans are rarely feasible due to the unpredictable nature of real application environments. Most software retains user preferences (e.g., remember the last configuration of user), meaning that when executing a specific task, the plan $p$ generated by the Planner might not align with the actual state of the software. Therefore, the model must determine whether to proceed with a subtask $S_{t}$ based on the current state (screenshot: $V_{t}$, metadata: $M_{t}$).
As illustrated in Figure 8, we employ an MLLM to determine whether the current task has been completed or requires modification. We systematically categorize the possible outcomes into:
(1) <Modify>: Indicates that the subtask should be modified or additional subtasks should be added.
(2) <Pass>: Indicates that the current subtask is unnecessary and can be skipped.
(3) <Continue>: Indicates that the subtask is valid and should be executed as planned.
(4) <Finished>: Indicates that the subtask has already been completed and requires no further action.
In cases where the screenshot does not provide sufficient visual information for the MLLM to determine the output, the model outputs #\#Cannot confirm”. When this occurs, we design a Region Search module implemented by an LLM. This module takes the corresponding GUI information extracted by the GUI parser and the task description of the current subtask to identify the relevant region. It then crops the region using the generated bounding box as the center coordinate, with the maximum width and height set to half of the original screenshot dimensions (ensure the region is smaller than the original screenshot). The cropped screenshot is subsequently sent to the Step-Check module to regenerate the decision.
B.4 Actor
The goal of the Actor is to translate natural language subtask $S_{t}$ into executable code $C_{t}$. Using an MLLM as the backbone model, the Actor processes metadata $m_{t}$ and screenshot $V_{t}$ as GUI context to generate precise executable actions, such as click(100, 200). Additionally, it leverages the history of previous actions as memory to aid in generating subsequent actions. The generated actions will be executed in the environment, and then the new screenshot $V_{t+1}$ and metadata $m_{t+1}$ will be captured for the next processing.
B.5 Actor-Critic
Figure 9: Actor-Critic. This module includes two parts: task verification and task correction. The design follows the verify-then-correct mechanism.
[…]
Table 11: Performance comparison between WorldGUI-Agent (with Claude-3.5-Sonnet and Claude-Sonnet-4) and Agent-S2 (Claude-Sonnet-4). Results are reported across five representative applications.
Method PPT VSCode Acrobat VLC File Explorer
Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
Agent-S2 w/ Claude-Sonnet-4 45.5 18.9 45.5 37.8 18.2 14.5 20.0 16.3 60.0 64.7
WorldGUI-Agent w/ Claude-3.5-Sonnet 54.5 39.6 54.5 26.7 63.6 20.0 50.0 18.4 50.0 17.6
WorldGUI-Agent w/ Claude-Sonnet-4 63.6 52.8 54.5 28.9 54.5 30.9 40.0 28.6 70.0 26.5
Table 12: Performance comparison between WorldGUI-Agent (with UI-TARS-1.5) and UI-TARS-1.5.
Method PPT Acrobat
Meta Aug. Meta Aug.
WorldGUI-Agent w/ UI-TARS-1.5 36.6 18.9 36.4 9.1
UI-TARS-1.5 27.3 17.0 9.1 1.8
Appendix C Additional Experiments
As shown in Table 11, Agent-S2 (Agashe et al., 2025b) shows competitive results as compared with our WorldGUI-Agent. We also test on two representative office software to compare the effectiveness of our proposed agentic framework by replacing the base model with UI-TARS-1.5 (Qin et al., 2025) in Table 12. It is noted that to improve the performance of UI-TARS-1.5, we use the GPT-4o to task planning, as we found that UI-TARS struggles with understanding complex desktop software layout and cannot capture the dynamic initial condition changes. We use GPT-4o for better implementation.
Appendix D Data
D.1 Annotators
In this work, we have four annotators: A, B, C, and D. The team comprises one PhD student, one Master’s student, and two undergraduate students. Prior to annotation, all annotators receive training on using the applications in WorldGUI to ensure high-quality annotations. For the 10 desktop applications, we divide the software into four parts, assigning each part to a different annotator. For the human tests presented in Table 4, the annotators demonstrate tasks on software that they did not annotate. As shown in Table 1, each annotator is responsible for different software during both the annotation and human testing phases to make the soundness of the Human Test results.
Table 13: The annotation arrangement during the annotation and human testing phases by different annotators.
Annotators Annotation Phase Human Test Phase
A PowerPoint, Word, Excel VSCode, VLC Player, Web
B Adobe Acrobat, VLC Player Excel, Settings
C Settings, Web PowerPoint, File Explorer, Youtube
D VSCode, File Explorer Word, Adobe Acrobat
Figure 10: Pipeline of Data Construction. Human: Represents the annotators. Code: Refers to the scripts (e.g., Python Code) utilized to achieve the goal. Agent: We design an agent built upon the MLLMs to achieve the goal.
D.2 Creating Augmented Tasks
In our study, to simulate dynamic testing processing in real GUI interactions, we propose to design GUI tasks with various initial tasks. Specifically, we propose pre-actions before executing the task. The pre-actions primarily serve two purposes: 1) Simulating Intermediate Task States: Pre-actions can complete specific steps of a task, creating a starting point from an intermediate state. This approach addresses scenarios where users may seek AI assistance because they are unable to complete a task. For example, if the task involves opening a dropdown menu, the pre-action may pre-open the menu. If the agent fails to recognize this precondition and follows its plan to click the menu again, it might inadvertently close the menu, causing task failure.
D.3 Introducing Diverse Initial Context States
Pre-actions can also introduce variations in the initial state, such as opening random tabs or settings. This ensures that the starting state is unconventional, challenging the agent to adapt by modifying its plan or adding new steps. We illustrate one example in Figure 5. Here, the meta task and augmented task, have the same user query and instructional video and it will ideally have the same final state. We additionally provide more examples about augmenting the meta task in Figure 11.
Figure 11: We present the examples of conducting the augmentations from the meta task.
Appendix E Detailed Experimental Results
Table 14 shows the detailed results of WorldGUI-Agent across individual applications in WindowsAgentArena (Bonatti et al., 2024) benchmark. The results of this related Windows-centric interactive GUI benchmark indicate that current the desktop GUI tasks are more challenging than web tasks. As we complete 11 out of 17 tasks in Web Browsing, a similar phenomenon is also discovered in Table 4.
Table 14: Detailed experimental results of WorldGUI-Agent across individual applications in WindowsAgentArena (Bonatti et al., 2024).
Domain Application #Tasks #Successes SR (%)
Web Browsing chrome 17 11 64.71
Windows Utilities clock 4 2 50.00
Windows System file_explorer 19 7 36.84
Office libreoffice_calc 24 1 4.17
Office libreoffice_writer 19 2 10.53
Windows Utilities microsoft_paint 3 1 33.33
Web Browsing msedge 13 5 38.46
Windows Utilities notepad 2 1 50.00
Windows System settings 5 4 80.00
Media & Video vlc 21 6 28.57
Coding vs_code 24 8 33.33
Windows Utilities windows_calc 3 0 0.00
Overall 154 48 31.17
Appendix F Computational Costs Discussion
The average number of execution steps and tokens consumed are shown below Table 15. The execution steps are calculated based on our experimental log files, while the token costs are sampled from representative tasks in each category by taking Actor module as an example.
Table 15: Average execution steps and token costs on different software.
Application category Average execution steps Input tokens per step (Actor) per task Output tokens per step (Actor) per task
Office $\sim$23 2350 212
Win. Usage $\sim$20 1929 108
Web $\sim$17 1637 84
Take a Windows Setting task as an example, we provide detailed time costs across different modules tested on a desktop PC with AMD Ryzen 7 5800H CPU. Task length: 6 (generated by Planner+Planner-Critic). To facilitate a fair comparison, we additionally selected two of the latest SOTA agents, Agent-S (2024-10-08) and Agent-S2 (2025-04-01), and measured their run times on the same successful task under identical hardware and the same base MLLM (Claude-Sonnet-4). The results are shown in Tables 16, 17, 18. To summarize, our WorldGUI-Agent shows a competitive running time of 129.55s, as compared with Agent-S (131.98s) and Agent-S2 (108.64s). The main computational costs of our designed modules are largely affected by the underlying large multimodal model, leaving room for acceleration optimization.
Since desktop GUI automation is still in its early stages, such computational costs are currently unavoidable. For reference, even OpenAI’s Deep Research reportedly takes over 10 minutes in daily usage. According to OpenAI Operator’s report, achieving 38.1% on OS-World requires over 100 steps, which is similarly costly. In summary, there remains a clear tradeoff between performance and time costs in GUI automation, and this challenge is shared across the community.
Table 16: Running time with WorldGUI-Agent (ours).
Subtask Index Executed Modules Time (seconds)
0 Planner 4.48
0 Planner-Critic 11.47
1 Parser 2.03
1 Step-Check 4.95
1 Actor 7.47
1 Parser 2.07
1 Actor-Critic 7.38
2 Parser 2.03
2 Step-Check 6.71
2 Actor 6.05
2 Parser 2.07
2 Actor-Critic 7.97
3 Parser 2.12
3 Step-Check 6.35
3 Actor 6.71
3 Parser 2.22
3 Actor-Critic 10.06
4 Parser 2.01
4 Step-Check 6.19
4 Actor 8.54
4 Parser 2.40
4 Actor-Critic 9.55
5 Parser 2.22
5 Step-Check 6.50
Total - 129.55
Average (per action) - 22.72
Table 17: Running time with Agent-S.
Step Executed Modules Time (seconds)
0 Manager 2.02
1 Manager 10.11
2 Worker 16.05
3 Worker 12.38
4 Worker 17.30
5 Worker 12.26
6 Worker 15.17
7 Worker 15.46
8 Worker 11.83
9 Worker 19.40
Total - 131.98
Average (per worker) - 14.98
Table 18: Running time with Agent-S2.
Step Executed Modules Time (seconds)
0 Manager 11.87
1 Worker 7.27
2 Worker 13.21
3 Worker 13.59
4 Worker 14.04
5 Manager 10.61
6 Worker 9.87
7 Manager 7.53
8 Worker 20.65
Total - 108.64
Average (per worker) - 13.22
Appendix G Examples of Augmentations
In this section, we present several augmentation examples in Figures 12, 13, 14, 15, 16, 17, 18. It is noted that our augmentations are not only making the first step changing but also require the agent add new step in its second step. For instance, in Figure 12, our augmentation is about click on Data tab in the ribbon, in the default software state, the Merge & Center button exhibit in the Home tab, there is no need to click on Home tab, after our augmentations, the agent should add a new task “Click on Home Tab” before it click on the Merge & Center button. Similarly, in Figure 17, the text editing buttons are under the Home Tab, if we augment the initial state with other Tab like Animation Tab, after the first step “Select the text ’US SUBMARINE DAY’ ”, the agent should add a new step like “Click on Home Tab” back to the default state for task execution. Except for adding new steps, we also present an example about adjust the step in Figure 13, as the target is about merging cells A1 to K1, we augment the initial state by selecting A2 to K2. Such a slight difference may mislead the agent to perceive such a minor difference, and the agent may jump the first step about selecting the correct cells lead to finally unsucess. In Figure 14 and Figure 18, we show two examples of introducing pop-up window in the initial state which require the agents accurately identify the pop-up windows and correctly close it by replanning the task based on the visual screenshot not only strictly planning based on inherited knowledge or the instructinal videos. In Figure 15, we show an example of changing the interface by clicking the Data tab to hide the Merge & Center button under the Home tab. In Figure 16, we complete the first step about selecting A1 to K1, which requires the agent to jump this step to reduce the time costs.
Figure 12: Augmented example of an Excel Task.
Figure 13: Augmented example of an Excel Task.
Figure 14: Augmented example of an Excel Task.
Figure 15: Augmented example of an Excel Task.
Figure 16: Augmented example of an Excel Task.
Figure 17: Augmented example of a PowerPoint Task.
Figure 18: Augmented example of a PowerPoint Task.
Appendix H WorldGUI-Agent Reasoning Loop Algorithm
In this section, we provide the details of our reasoning loop algorithm in Algorithm 1.
Algorithm 1 WorldGUI-Agent Reasoning Loop Algorithm
Input: State $s$, Action Code $C$, Screenshot $V$, Metadata $m$, Current subtask $S$, Critic_count $z$
 Generate task plan $p$ with Planner and Planner-Critic
 Initial current subtask $S_{t=0}=S_{1}^{1}$, where $S_{1}^{0}$ is the $1$-th subtask in the $1$-th milestone of pp.
 Initial $s_{0}=\textless Continue\textgreater$
while $S_{t}$ is not end and $t<$ max trials do
  Observe metadata $m_{t}$ and Screenshot $V_{t}$ from Env.
  Obtain state $s_{t}$ by running Step-Check.
  if $s_{t}=\textless Next\textgreater$ then
   Go to the next task $S_{t+1}=next(S_{t})$
  end if
  Check potential modification of subtask $S_{t}$
  Obtain action code $C_{t}$ by running Actor; Execute the action code $C_{t}$ in the Env.; Observe metadata $m_{t}$ and Screenshot $V_{t}$ from Env.
  Set $C_{t}=$ None; $t=t+1$; Set state $s_{t}=\textless Critic\textgreater$ (For each subtask, the first step is finished, then execute the actor-critic process)
  Observe metadata $m_{t}$ and Screenshot $V_{t}$ from Env.
  Running Actor-Critic and obtain the state $s_{t}$
  if $s_{t}=\textless Next\textgreater$ then
   Go to the next task $S_{t+1}=next(S_{t})$.
  end if
  while $s_{t}=\textless Critic\textgreater$ and $z<$ max critique trials do
   Running Actor-Critic and obtain the state $s_{t}$ and corrected action code $C_{t}$
   if $s_{t}=\textless Next\textgreater$ then
    Go to the next task $S_{t+1}=next(S_{t})$.
   end if
   Execute the action code $C_{t}$ in the Env.; Observe metadata $m_{t}$ and Screenshot $V_{t}$ from Env.
   Set $C_{t}=$ None; $z=z+1$
  end while
  Go to the next task $S_{t+1}=next(S_{t})$
  $t=t+1$
end while
Appendix I Qualitative Results
(1) In Figure 19, we present a successful prediction example, demonstrating that the WorldGUI can effectively plan each step for a task, accurately perceive specific elements in the GUI, and convert them into the correct action code. Additionally, we display the parsed GUI elements, which can accurately identify most content, including small icons and dense text elements. (2) We provide the visualization results of using Planner-Critic, Step-Check, and Actor-Critic in Figure 21, Figure 22, and Figure 23. These qualitative results demonstrate the effectiveness of these critical modules in GUI automation. (3) We also highlight some common errors encountered. 1) The model has difficulty obtaining the desired information when we augment the task by invoking the dropdown menu of the Settings application. As shown on the left of Figure 24, when we click on the ’System’ button on the left, it is challenging for our model to extract the button’s position as it is hidden. Such cases require the model to have a higher level of ability to delete the content in the input box or click on the blank area. 2) As shown in the right of Figure 24, the model has difficulty dragging a bar to achieve the desired value. 3) The model struggles with the visual choice when there is no text information in the screenshot, as shown on the left of Figure 25. The subtask aims to select the center button, but the current model makes it hard to detect the center choice only from the screenshot. 4) The model cannot successfully locate the position of the input box, as the GUI parser will easily locate the text location ’Replace with’, it always outputs the action like clicking on the ’Replace with’, which will destroy the whole task’s success.
Figure 19: We show one successful prediction of our WorldGUI-Agent.
Figure 20: We show two examples of using GUI Parser to obtain the element position information.
Figure 21: An example of using Planner-Critic to correct the plan.
Figure 22: Two examples of using Step-Check to check the subtask status.
Figure 23: An example of using Actor-Critic to correct the actions.
Figure 24: We display some common errors.
Figure 25: We display some common errors